# Let Me Speak Freely? A Study on the Impact of Format Restrictions on Performance of Large Language Models

### v1

@@ Abstract
Structured generation, the process of producing content in standardized formats like JSON and XML, is widely utilized in real-world applications to extract key output information from large language models (LLMs). This study investigates whether such constraints on generation space impact LLMs’ abilities, including reasoning and domain knowledge comprehension. Specifically, we evaluate LLMs’ performance when restricted to adhere to structured formats versus generating free-form responses across various common tasks. Surprisingly, we observe a significant decline in LLMs’ reasoning abilities under format restrictions. Furthermore, we find that stricter format constraints generally lead to greater performance degradation in reasoning tasks.
@@ 1 Introduction
The few-shot in-context learning Brown et al. (2020) and instruction-following Wei et al. (2021) capabilities of large language models (LLMs) have enabled their out-of-the-box usage to solve downstream tasks. However, a major obstacle to incorporating LLMs into industrial applications is their lack of adherence to standardized output formats. This inconsistency complicates output parsing and undermines the reliability of these models.
@@ 3.3 Evaluation method
Consideration for Prompt Sensitivity. Previous studies (Chen et al., 2023; Sclar et al., 2023; Zhu et al., 2023) have shown that LLMs are sensitive to slight variations in prompts. To account for this, we evaluate our approach by nine prompt combinations: three task descriptions and three JSON, XML, and YAML schemas with slight variations in wording or format. For natural language prompting, we include three variations in text formats (e.g., Give your reason first followed by your answers). Details of the task description prompts and FRI prompts can be found in Appendix F.
@@ 5.2 Comparison Across Different Formats
Figure 4: Comparison of different format in classification related tasks on DDXPlus, Sports, Task280 and Multifin. NL=Natural Language. We showed the averaged accuracy for each format over 9 different prompts with standard deviation error.
In this section we ablate the format language by comparing not just JSON but also XML and YAML format. Since all 3 language comes in different grammar syntax rules and restriction. We deduce each models might perform differently for example Claude-3-Haiku uses XML for tool use schema so
On hint sight we do not see any structure format which consistency stands out from others which generalized across all models in Figure 4. For Gemini model, we found JSON is more consistent however it does not always outperform other format.
@@ 5.3 Structure Format and Parsing Error Rates
We initially hypothesized that the performance gap between text and structured formats might be attributed to parsing errors during answer extraction. However, our analysis of error rates across different formats and models, as shown in Table 2, reveals that this is not the primary factor. In fact, Gemini 1.5 Flash and GPT 3.5 Turbo exhibit near zero parsing failures in all three format. In the LLaMA 3 8B setting, the parsing error rate for the Last Letter task in JSON format is only 0.148%, yet there exists a substantial 38.15% performance gap as seen in Table 1.
@@ 6 Related Work
The exploration of LLMs’ ability to follow instructions and produce responses in specified formats was first addressed by IFEval Zhou et al. (2023) which designed to evaluate the general instruction-following ability of LLMs, and it contains a subset of test instances specifically assessing format-following. INFOBENCH Qin et al. (2024) introduces a broader coverage of instructions and conducts a more fine-grained analysis by decomposing the instructions into different categories, including format specifications. FOFO Xia et al. (2024) is a benchmark solely focused on the format-following ability of LLMs. However, these works do not explore if format instruction interfere with downstream performance.
@@ 7 Conclusion
Our study reveals that structured generation constraints significantly impact LLM performance across various tasks. Format restrictions, particularly constrained decoding (JSON-mode), can hinder reasoning abilities while enhance classification task accuracy. Looser format restrictions generally improve performance and reduce variance in reasoning tasks. Parsing errors, while not the primary cause of performance differences, can be mitigated through corrective prompting. These findings underscore the importance of balancing format adherence, reasoning capabilities, and cost efficiency in LLM applications. Given that our study focuses on reasoning-intensive tasks, future work should explore how reasoning tasks of varying difficulty, from intensive to simple, are affected by restrictive formats and LLMs. To mitigate the performance degradation of LLMs due to restrictive formats, future studies should include a wider range of training data that contains instructions in various restrictive formats in local LLMs.
@@ Appendix E Comparison between using regex and LLM as answer parser in GSM8K
To answer the difference between using regex parser to extract the final strict match answer, we calculate the Exact Match score in GSM8K results using the prompt format template "The final answer is". Table 5 results reveal a significant gap between regex match and LLM as final answer parser in EM score across various language models, highlighting the limitations of using only one strict regex matching for different models. For example, GPT-3.5-Turbo shows a 31.8 percentage point improvement from regex match (43.7%) to overall accuracy (75.5%), while Gemini-1.5-Flash exhibits an even larger 43.5 point difference. This pattern is consistent across all models, with mistral-7b demonstrating the most dramatic 42 point increase. These disparities underscore the value of using LLMs as answer parsers, as they can understand and evaluate responses beyond literal string matching, accounting for paraphrases and contextual understanding, thus providing a more nuanced and accurate assessment in text-based tasks.
@@
Appendix F Prompt
F.1 Prompt Format
@@
Format Description A format description includes the target format (ie JSON, XML or YAML) and a targeted schema we intent the LLM response to adhere to.
@@
F.2 Prompt variation for GSM8K
The 3 task description variations for GSM8K are shown in as below:
@@ F.2 Prompt variation for GSM8K
Format descriptions are shown in Section F.10.
F.3 Prompt variation for Last Letter Concatenation
The 3 task description variations for Last Letter Concatenation are shown in as below:
@@ F.3 Prompt variation for Last Letter Concatenation
F.4 Prompt variation for Shuffle Object Tracking
The 3 task description variations for Shuffle Object Tracking are shown in as below:
@@ F.4 Prompt variation for Shuffle Object Tracking
F.5 Prompt variation for DDXPlus
The 3 task description variations for DDXPlus are shown in as below: {mdframed} Task description variation1:
@@ F.5 Prompt variation for DDXPlus
- Possible NSTEMI / STEMI - Spontaneous rib fracture - Pulmonary embolism - Pulmonary neoplasm - Scombroid food poisoning RESPONSE: Task description variation2:
@@
All possible valid diagnoses for you to choose from are as follows: - Possible NSTEMI / STEMI - Spontaneous rib fracture - Pulmonary embolism - Pulmonary neoplasm - Scombroid food poisoning Task description variation3:
@@
• Given: A patient profile • Required: Diagnose the patient based on the provided information • Process: Think step by step to analyze the patient’s symptoms and history • Output: Select one diagnosis from the provided list of valid options Note: Carefully review the patient profile and the list of possible diagnoses before making your determination. Do not answer "Insufficient information" - you must choose from the given options. Valid diagnoses (select one): - Possible NSTEMI / STEMI - Spontaneous rib fracture - Pulmonary embolism - Pulmonary neoplasm - Scombroid food poisoning
F.6 Prompt variation for Sports Understanding
The 3 task description variations for Sports Understanding are shown in as below: {mdframed} Task description variation1:
@@ F.6 Prompt variation for Sports Understanding
Read carefully for each of the last question and think step by step before answering. Answer yes if its plausible, no if implausible Task description variation2:
@@
• Given: A sentence related to sports • Required: Determine if the sentence is plausible or implausible • Process: Think step by step to analyze the sentence • Output: Answer "yes" if plausible, "no" if implausible
F.7 Prompt variation for Natural Language - Task 280
The 3 task description variations for Natural Language - Task 280 are shown in as below:
@@ F.7 Prompt variation for Natural Language - Task 280
F.8 Prompt variation for MultiFin
The 3 task description variations for MultiFin are shown in as below:
@@ F.8 Prompt variation for MultiFin
F.9 Direct Answer format prompts
For tasks which doesn’t need chain of though reasoning, we use these three prompt variations. In our experiments, we use 3 format variations in MultiFin.
@@ F.9 Direct Answer format prompts
F.10 Chain of thought format prompts
For tasks which include chain of thought reasoning, we use these three prompt variations. In our experiments, we use 3 format variations in GSM8K, Last Letter, Shuffle Object Tracking, DDXPlus, Sports, NL-Task 280.
@@ F.10 Chain of thought format prompts
Appendix G Averaged numbers for all datasets
G.1 Zero shot prompting comparing Text, JSON, XML, YAML
Table LABEL:tab:number_breakdown shows all the number with standard deviation on all 4 format (NL, JSON, XML, YAML).
The JSON-mode scores for GPT 3.5 turbo, Gemini 1.5 Flash and LLaMA 3 8B
Last Letter GSM8K ShuffleObj
Gemini-1.5-Flash
   Text 65.45 (3.1) 89.33 (0.8) 58.21 (13.0)
   JSON 77.02 (7.3) 89.21 (1.5) 65.07 (5.3)
   XML 74.17 (10.4) 88.20 (2.2) 50.39 (10.5)
   YAML 71.43 (20.3) 87.42 (3.7) 34.27 (17.1)
GPT-3.5 Turbo
   Text 56.74 (7.1) 76.60 (2.8) 20.37 (3.6)
   JSON 25.20 (29.1) 49.25 (12.0) 20.93 (1.1)
   XML 22.30 (27.8) 45.06 (19.9) 15.35 (1.8)
   YAML 66.87 (22.0) 73.85 (5.6) 20.79 (1.3)
Claude 3 Haiku
   Text 57.67 (21.1) 86.51 (0.8) 36.62 (8.2)
   JSON 56.74 (16.7) 23.44 (22.8) 49.33 (4.8)
   XML 33.80 (31.5) 79.76 (7.0) 48.10 (5.2)
   YAML 31.60 (32.4) 80.63 (2.8) 18.10 (14.7)
LLaMA 3 8B
   Text 70.07 (5.3) 74.73 (0.6) 27.01 (5.5)
   JSON 28.00 (12.2) 48.90 (6.7) 15.72 (11.0)
   XML 15.93 (4.8) 56.74 (8.3) 11.14 (3.6)
   YAML 16.10 (10.4) 46.08 (16.8) 9.57 (3.6)
Gemma2 9B IT
   Text 56.81 (9.8) 86.50 (0.6) 49.39 (5.8)
   JSON 38.96 (6.8) 84.17 (3.7) 50.49 (8.9)
   XML 43.70 (3.8) 85.56 (0.6) 38.47 (5.0)
   YAML 23.41 (15.7) 79.52 (4.1) 22.96 (16.4)
DDXPlus Sports Task280 MultiFin
Gemini-1.5-Flash
   Text 41.59 (6.6) 79.91 (3.2) 68.60 (2.5) 63.45 (0.3)
   JSON 60.36 (0.8) 78.85 (1.3) 70.27 (5.4) 65.24 (1.1)
   XML 59.44 (1.4) 80.15 (0.7) 69.96 (4.9) 64.54 (1.6)
   YAML 60.41 (1.6) 80.10 (1.2) 65.31 (12.7) 64.14 (0.4)
GPT-3.5 Turbo
   Text 44.07 (3.2) 67.15 (26.8) 72.65 (6.3) 63.03 (0.5)
   JSON 55.51 (0.4) 79.99 (3.3) 70.56 (11.2) 64.03 (0.9)
   XML 52.98 (1.4) 80.71 (1.1) 66.22 (16.2) 62.24 (1.1)
   YAML 55.01 (0.8) 80.93 (2.3) 72.10 (8.0) 65.38 (0.9)
Claude 3 Haiku
   Text 33.78 (13.5) 77.81 (5.8) 61.09 (11.0) 61.95 (1.9)
   JSON 52.04 (1.1) 78.74 (2.8) 49.46 (27.2) 63.68 (1.3)
   XML 50.75 (0.8) 77.78 (3.8) 45.03 (25.0) 62.39 (1.1)
   YAML 6.91 (5.3) 76.37 (8.3) 44.51 (24.2) 61.75 (1.7)
LLaMA 3 8B
   Text 12.04 (15.2) 69.49 (12.7) 65.28 (3.4) 60.26 (1.4)
   JSON 23.37 (0.7) 73.38 (3.5) 39.46 (22.4) 57.74 (2.0)
   XML 11.35 (1.9) 69.20 (5.5) 35.36 (22.5) 58.77 (3.2)
   YAML 13.08 (4.1) 68.25 (5.7) 45.42 (24.4) 49.74 (4.2)
Gemma2 9B IT
   Text 22.86 (5.8) 76.08 (2.3) 69.84 (7.7) 69.69 (0.4)
   JSON 52.97 (0.2) 72.72 (1.6) 65.57 (11.7) 70.18 (0.7)
   XML 52.92 (2.8) 73.29 (2.4) 68.07 (11.7) 67.98 (0.7)
   YAML 44.94 (2.2) 72.97 (1.7) 60.45 (11.0) 69.81 (0.7)
Dataset GPT3.5T Flash LLaMA3
LastLetter 1.78 (0.3) 0.67 (0.5) 7.56 (2.7)
GSM8K 29.87 (0.8) 47.78 (3.1) 65.38 (1.3)
MultiFin 66.00 (1.3) 66.79 (0.4) 54.82 (1.5)
Sports 76.82 (0.9) 77.79 (0.4) 72.08 (2.6)
Task 280 78.07 (2.3) 67.19 (4.1) 74.57 (2.0)
DDXPlus 51.87 (2.8) 84.92 (2.1) 22.59 (0.1)
Table 8: Applying JSON-mode to all 6 datasets and found

### v2

@@ Abstract
Structured generation, the process of producing content in standardized formats like JSON and XML, is widely utilized in real-world applications to extract key output information from large language models (LLMs). This study investigates whether such constraints on generation space impact LLMs’ abilities, including reasoning and domain knowledge comprehension. Specifically, we evaluate LLMs’ performance when restricted to adhere to structured formats versus generating free-form responses across various common tasks. Surprisingly, we observe a significant decline in LLMs’ reasoning abilities under format restrictions. Furthermore, we find that stricter format constraints generally lead to greater performance degradation in reasoning tasks. Our code and results are available online. [^1] https://github.com/appier-research/structure-gen
@@ 1 Introduction
The few-shot in-context learning Brown et al. (2020) and instruction-following Wei et al. (2021) capabilities of large language models (LLMs) have allowed them to solve downstream tasks out of the box. However, a major obstacle to incorporating LLMs into industrial applications is their lack of adherence to standardized output formats. This inconsistency complicates output parsing and undermines the reliability of these models.
@@ 3.3 Evaluation method
Consideration for Prompt Sensitivity. Previous studies (Chen et al., 2023; Sclar et al., 2023; Zhu et al., 2023; Mizrahi et al., 2024) have shown that LLMs are sensitive to slight variations in prompts. To account for this, we evaluate our approach by nine prompt combinations: three task descriptions and three JSON, XML, and YAML schemas with slight variations in wording or format. For natural language prompting, we include three variations in text formats (e.g., Give your reason first followed by your answers). Details of the task description prompts and FRI prompts can be found in Appendix G.
@@ 5.2 Comparison Across Different Formats
Figure 4: Comparison of different formats in classification related tasks on DDXPlus, Sports, Task280 and Multifin. NL=Natural Language. We showed the averaged accuracy for each format over 9 different prompts with standard deviation error.
In this section we ablate the format language by comparing not just JSON but also XML and YAML format. Since all 3 language comes in different grammar syntax rules and restriction. We expect each models might perform differently for example Claude-3-Haiku uses XML for tool use schema.
On hindsight we do not see any structure format which consistency stands out from others which generalized across all models in Figure 4. For Gemini model, we found JSON is more consistent however it does not always outperform other format.
@@ 5.3 Structure Format and Parsing Error Rates
We initially hypothesized that the performance gap between text and structured formats might be attributed to parsing errors during answer extraction. However, our analysis of error rates across different formats and models, as shown in Table 2, reveals that this is not the primary factor. In fact, Gemini 1.5 Flash and GPT 3.5 Turbo exhibit near zero parsing failures in all three formats. In the LLaMA 3 8B setting, the parsing error rate for the Last Letter task in JSON format is only 0.148%, yet there exists a substantial 38.15% performance gap as seen in Table 1.
@@ 6 Related Work
The exploration of LLMs’ ability to follow instructions and produce responses in specified formats was first addressed by IFEval Zhou et al. (2023) which aimed to evaluate the general instruction-following ability of LLMs, and it contains a subset of test instances specifically assessing format-following. INFOBENCH Qin et al. (2024) introduces a broader coverage of instructions and conducts a more fine-grained analysis by decomposing the instructions into different categories, including format specifications. FOFO Xia et al. (2024) is a benchmark solely focused on the format-following ability of LLMs. However, these works do not explore if format instruction interfere with downstream performance.
@@ 7 Conclusion
Our study reveals that structured generation constraints significantly impact LLM performance across various tasks. Format restrictions, particularly constrained decoding (JSON-mode), can hinder reasoning abilities while enhancing classification task accuracy. Looser format restrictions generally improve performance and reduce variance in reasoning tasks. Parsing errors, while not the primary cause of performance differences, can be mitigated through corrective prompting. These findings underscore the importance of balancing format adherence, reasoning capabilities, and cost efficiency in LLM applications. Given that our study focuses on reasoning-intensive tasks, future work should explore how reasoning tasks of varying difficulty, from intensive to simple, are affected by restrictive formats and LLMs. To mitigate the performance degradation of LLMs due to restrictive formats, future studies should include a wider range of training data that contains instructions in various restrictive formats in local LLMs.
@@ Appendix E Comparison between using regex and LLM as answer parser in GSM8K
To answer the difference between using regex parser to extract the final strict match answer, we calculate the Exact Match score in GSM8K results using the prompt format template "The final answer is". Table 6 results reveal a significant gap between regex match and LLM as final answer parser in EM score across various language models, highlighting the limitations of using only one strict regex matching for different models. For example, GPT-3.5-Turbo shows a 31.8 percentage point improvement from regex match (43.7%) to overall accuracy (75.5%), while Gemini-1.5-Flash exhibits an even larger 43.5 point difference. This pattern is consistent across all models, with mistral-7b demonstrating the most dramatic 42 point increase.
These disparities underscore the value of using LLMs as answer parsers, as they can understand and evaluate responses beyond literal string matching, accounting for paraphrases and contextual understanding, thus providing a more nuanced and accurate assessment in text-based tasks.
Just to be safe we also assess the reliability of GPT-4-turbo as a parser, we conducted a manual validation study:
  • We sampled 300 responses in total: 100 each from Last Letter, Shuffle Object, and GSM8K
  • These responses were independently parsed by human evaluators.
  • We then compared the human-parsed answers with those extracted by GPT-4-turbo.
The results of this validation are shown in Table 5. These findings demonstrate an average alignment of 97.7% between GPT-4-turbo and human-parsed answers, supporting our characterization of GPT-4-turbo as a near-perfect parser for this task.
Task GPT-4-Turbo correctness
Last Letter 97/100
Shuffle Obj 96/100
GSM8K 100/100
Table 5: Alignment between GPT-4-turbo and human-parsed answers
@@
Appendix F Averaged numbers for all datasets
F.1 Zero shot prompting comparing Text, JSON, XML, YAML
Table (8, 7) shows all the number with standard deviation on all 4 format (NL, JSON, XML, YAML) in classification and reasoning tasks.
The JSON-mode scores for GPT 3.5 turbo, Gemini 1.5 Flash and LLaMA 3 8B are presented in Table 9. This table shows the performance of these three models on six different datasets when using JSON-mode.
Last Letter GSM8K ShuffleObj
Gemini-1.5-Flash
   Text 65.4 (3.1) 89.3 (0.8) 58.2 (13.0)
   JSON 77.0 (7.3) 89.2 (1.5) 65.1 (5.3)
   XML 74.2 (10.4) 88.2 (2.2) 50.4 (10.5)
   YAML 71.4 (20.3) 87.4 (3.7) 34.3 (17.1)
GPT-3.5 Turbo
   Text 56.7 (7.1) 76.6 (2.8) 20.4 (3.6)
   JSON 25.2 (29.1) 49.3 (12.0) 20.9 (1.1)
   XML 22.3 (27.8) 45.1 (19.9) 15.4 (1.8)
   YAML 66.9 (22.0) 73.9 (5.6) 20.8 (1.3)
Claude 3 Haiku
   Text 57.7 (21.1) 86.5 (0.8) 36.6 (8.2)
   JSON 56.7 (16.7) 23.4 (22.8) 49.3 (4.8)
   XML 33.8 (31.5) 79.8 (7.0) 48.1 (5.2)
   YAML 31.6 (32.4) 80.6 (2.8) 18.1 (14.7)
LLaMA 3 8B
   Text 70.1 (5.3) 74.7 (0.6) 27.0 (5.5)
   JSON 28.0 (12.2) 48.9 (6.7) 15.7 (11.0)
   XML 15.9 (4.8) 56.7 (8.3) 11.1 (3.6)
   YAML 16.1 (10.4) 46.1 (16.8) 9.6 (3.6)
Gemma2 9B IT
   Text 56.8 (9.8) 86.5 (0.6) 49.4 (5.8)
   JSON 39.0 (6.8) 84.2 (3.7) 50.5 (8.9)
   XML 43.7 (3.8) 85.6 (0.6) 38.5 (5.0)
   YAML 23.4 (15.7) 79.5 (4.1) 23.0 (16.4)
DDXPlus Sports Task280 MultiFin
Gemini-1.5-Flash
   Text 41.6 (6.6) 79.9 (3.2) 68.6 (2.5) 63.5 (0.3)
   JSON 60.3 (0.8) 78.9 (1.3) 70.3 (5.4) 65.2 (1.1)
   XML 59.4 (1.4) 80.2 (0.7) 70.0 (4.9) 64.5 (1.6)
   YAML 60.4 (1.6) 80.1 (1.2) 65.3 (12.7) 64.1 (0.4)
GPT-3.5 Turbo
   Text 44.1 (3.2) 67.2 (26.8) 72.7 (6.3) 63.0 (0.5)
   JSON 55.5 (0.4) 80.0 (3.3) 70.6 (11.2) 64.0 (0.9)
   XML 53.0 (1.4) 80.7 (1.1) 66.2 (16.2) 62.2 (1.1)
   YAML 55.0 (0.8) 80.9 (2.3) 72.1 (8.0) 65.4 (0.9)
Claude 3 Haiku
   Text 33.8 (13.5) 77.8 (5.8) 61.1 (11.0) 62.0 (1.9)
   JSON 52.0 (1.1) 78.7 (2.8) 49.5 (27.2) 63.7 (1.3)
   XML 50.8 (0.8) 77.8 (3.8) 45.0 (25.0) 62.4 (1.1)
   YAML 6.9 (5.3) 76.4 (8.3) 44.5 (24.2) 61.8 (1.7)
LLaMA 3 8B
   Text 12.04 (15.2) 69.49 (12.7) 65.28 (3.4) 60.26 (1.4)
   JSON 23.37 (0.7) 73.38 (3.5) 39.46 (22.4) 57.74 (2.0)
   XML 11.35 (1.9) 69.20 (5.5) 35.36 (22.5) 58.77 (3.2)
   YAML 13.08 (4.1) 68.25 (5.7) 45.42 (24.4) 49.74 (4.2)
Gemma2 9B IT
   Text 22.9 (5.8) 76.1 (2.3) 69.8 (7.7) 70.0 (0.4)
   JSON 53.0 (0.2) 72.7 (1.6) 65.6 (11.7) 70.2 (0.7)
   XML 52.9 (2.8) 73.3 (2.4) 68.1 (11.7) 68.0 (0.7)
   YAML 44.9 (2.2) 73.0 (1.7) 60.5 (11.0) 69.8 (0.7)
Dataset GPT3.5T Gemini1.5F LLaMA3 8B
LastLetter 1.78 (0.3) 0.67 (0.5) 7.56 (2.7)
GSM8K 29.87 (0.8) 47.78 (3.1) 65.38 (1.3)
MultiFin 66.00 (1.3) 66.79 (0.4) 54.82 (1.5)
Sports 76.82 (0.9) 77.79 (0.4) 72.08 (2.6)
Task 280 78.07 (2.3) 67.19 (4.1) 74.57 (2.0)
DDXPlus 51.87 (2.8) 84.92 (2.1) 22.59 (0.1)
Table 9: Averaged scores for JSON-mode to all 6 datasets, performance varies significantly across tasks and models, suggesting that different models may have strengths in different areas when using JSON-mode.
Appendix G Prompt
G.1 Prompt Format
@@
Format Description A format description includes the target format (ie JSON, XML or YAML) and a targeted schema we intend the LLM response to adhere to.
@@
G.2 Prompt Variations
Our study employs a range of prompt variations across multiple tasks to assess the robustness and generalizability of language models. We developed three distinct task description variations for each of the following datasets:
  • GSM8K (Figure 9)
  • Last Letter (Figure 10)
  • Shuffle Object (Figure 11)
  • DDXPlus (Figure 12)
  • Sports Understanding (Figure 13)
  • Natural Language - Task 280 (Figure 14)
  • MultiFin (Figure 15)
For tasks involving chain-of-thought reasoning (GSM8K, Last Letter, Shuffle Object Tracking, DDXPlus, Sports Understanding, and NL-Task 280), we implemented three prompt format variations. These are illustrated in Figures 19, 20, and 21.
Additionally, we created three answering format variations for both reasoning-based tasks and those requiring direct answers. These "direct answer prompts" are presented in Figures 16, 17, and 18.
@@ G.2 Prompt Variations
Figure 9: GSM8K Task Description Variations
@@
Figure 10: Last Letter Task Description Variations
@@
Figure 11: Shuffle object Task Description Variations
{mdframed}
Task description variation1:
@@
- Possible NSTEMI / STEMI
- Spontaneous rib fracture
- Pulmonary embolism
- Pulmonary neoplasm
- Scombroid food poisoning
RESPONSE:
Task description variation2:
@@
All possible valid diagnoses for you to choose from are as follows:
- Possible NSTEMI / STEMI
- Spontaneous rib fracture
- Pulmonary embolism
- Pulmonary neoplasm
- Scombroid food poisoning
Task description variation3:
@@
• Given: A patient profile
• Required: Diagnose the patient based on the provided information
• Process: Think step by step to analyze the patient’s symptoms and history
• Output: Select one diagnosis from the provided list of valid options
Note: Carefully review the patient profile and the list of possible diagnoses before making your determination. Do not answer "Insufficient information" - you must choose from the given options.
Valid diagnoses (select one):
- Possible NSTEMI / STEMI
- Spontaneous rib fracture
- Pulmonary embolism
- Pulmonary neoplasm
- Scombroid food poisoning
Figure 12: DDXPlus Task Description Variations
{mdframed}
Task description variation1:
@@
Answer yes if its plausible, no if implausible
Task description variation2:
@@
• Given: A sentence related to sports
• Required: Determine if the sentence is plausible or implausible
• Process: Think step by step to analyze the sentence
• Output: Answer "yes" if plausible, "no" if implausible
Figure 13: Sports Task Description Variations
@@
Figure 14: Task 280 Task Description Variations
@@
Figure 15: MultiFin Task Description Variations
@@
Figure 16: Variation 1 for direct Answering format with only answer field in all 4 format.
@@
Figure 17: Variation 2 for direct Answering format with only answer field in all 4 format.
@@
Figure 18: Variation 3 for direct Answering format with only answer field in all 4 format.
@@
Figure 21: Reasoning response prompt - Variation 3